OpenStudy (anonymous):

List all real values of x such that f(x)=0. f(x)=6x^2-17x+9

OpenStudy (anonymous):

let's use the quadratic formula to solve \[\large 6x^2-17x+9=0 \] \[\large ax^2+bx+c=0 \] identify your a= , b= , c= and plug them into the quadratic formula: \[\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

yw... do you want an answer verification?

OpenStudy (anonymous):

\[\large x=\frac{17\pm\sqrt{73}}{12} \]